OpenStudy (anonymous):

Please help!! A wave is modeled with the function y = 1/2 sin 3 theta. Describe the graph of this function, including its period, amplitude, and points of intersection with the x-axis.

OpenStudy (campbell_st):

well the standard form of the curve is \[y = asin(bx)\] a = amplitude... the period is \[\frac{2\pi}{b}\] so comparing this to your equation what do you think is the amplitude and the period...?

OpenStudy (anonymous):

Amplitude: 1/2 Period: 2pi/3 Intersections: Infinitely many, but pi/3, 2pi/3, so on and so forth

OpenStudy (anonymous):

is tht correct

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

Thank you so much for your help and for explaining the process to me, many people jsut give me the answer without an explanation

OpenStudy (campbell_st):

glad to help